OpenStudy (anonymous):

The limit of (cos(pi*x) +1)/(x^2 -2x+1) as x tends to zero.

OpenStudy (agreene):

\[\lim_{x\rightarrow0}\frac{\cos(\pi x)+1}{x^2-2x+1}=2\]